Joe348:

the length of a rectangle is 4 meters less than twice its width. The area of the rectangle is 70m2. Find the dimensions of the rectangle

freshavacado850:

off topic but your pfp is hot

5lim3Ty:

@freshavacado850 wrote:
off topic but your pfp is hot
lmao

Bxmp:

Area=Length x Width. Length is 2w-4. So, \[70=w(2w-4)\] \[70=2w^2-4w\] \[2w^2-4w-70=0\] Factor the equation: \[2(w-7)(w+5)=0\] Can you continue from here?

Joe348:

2w-14 times w+5?

Bxmp:

No, \[w-7=0\] Add 7 to each side. \[w=7\] So, correct, it is 7 your width.

freshavacado850:

If the width = x m then the length = 2x - 4 m Area = x(2x - 4) = 70 2x^2 - 4x - 70 = 0 2x^2 - 14x + 10x - 70 = 0 2x(x - 7) + 10(x - 7) = 0 (2x + 10)(x - 7) = 0 x = 7 or -5 ( ignore negative) width = 7 m and length = 10 m answer

Bxmp:

Are you following so far?

Bxmp:

@freshavacado850 wrote:
If the width = x m then the length = 2x - 4 m Area = x(2x - 4) = 70 2x^2 - 4x - 70 = 0 2x^2 - 14x + 10x - 70 = 0 *2x(x - 7) + 10(x - 7) = 0* Wrong. (2x + 10)(x - 7) = 0 x = 7 or -5 ( ignore negative) width = 7 m and length = 10 m answer

Joe348:

trying to but not understanding

Joe348:

brb

Bxmp:

@bxmp wrote:
@freshavacado850 wrote:
If the width = x m then the length = 2x - 4 m Area = x(2x - 4) = 70 2x^2 - 4x - 70 = 0 2x^2 - 14x + 10x - 70 = 0 *2x(x - 7) + 10(x - 7) = 0* Wrong. (2x + 10)(x - 7) = 0 x = 7 or -5 ( ignore negative) width = 7 m and length = 10 m answer
And how did you get 10?, never showed any work for that.

Joe348:

where you get -7 from though?

Bxmp:

It's a positive seven, and from right here.

@bxmp wrote:
No, \[w-7=0\] Add 7 to each side. \[w=7\] So, correct, it is 7 your width.

freshavacado850:

@freshavacado850 wrote:
If the width = x m then the length = 2x - 4 m Area = x(2x - 4) = 70 2x^2 - 4x - 70 = 0 2x^2 - 14x + 10x - 70 = 0 2x(x - 7) + 10(x - 7) = 0 (2x + 10)(x - 7) = 0 x = 7 or -5 ( ignore negative) width = 7 m and length = 10 m answer
I got this from google also

Bxmp:

\[(w-7)\rightarrow(w-7=0)\rightarrow(w=7)\]

Bxmp:

Do you know how to get x-intercepts from the quadratic formula?